Using laws of exponents, simplify and write the answer in exponential form: 8^t ÷ 8^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of dividing the powers with same base = a^m÷ a^n= a^m - n
Then,
= (8)^t - 2

Thanks